OpenStudy (anonymous):

find the roots for f(x) = x4 + 21x2 – 100.

OpenStudy (anonymous):

@kc_kennylau do u know how to do this?

OpenStudy (raden):

first, setting f(x) = 0, in other words x^4 + 21x^2 - 100 = 0 to make be easier, let k = a^2. so the equation equals k^2 + 21k - 100 = 0 solve for k, first. like to get the roots of quadratic

OpenStudy (anonymous):

how do i solve for k ? @RadEn

OpenStudy (raden):

k^2 + 21k - 100 = 0 the equation above can be factored : (k + 25)(k - 4) = 0 take each factor equals zero, then solve for k

OpenStudy (raden):

k+25=0 k = .... ? also k-4=0 k= .... ?

OpenStudy (anonymous):

-25 & 4

OpenStudy (raden):

that's correct! now return k = a^2 so, we have 2 cases : a^2 = -25 and a^2 = 4 solve for a

OpenStudy (anonymous):

would it be 5, -5 and 2, -2?

OpenStudy (raden):

2 and -2, these roots are correct. but not yet for 5 and -5 a^2 = -25 a = +- sqrt(-25) a = +- sqrt(25 * -1) a = +- sqrt(25) sqrt(-1) a = +- 5 sqrt(-1) a = +- 5i i assumed the complex roots is allowed here

OpenStudy (anonymous):

so it would be -5i and 5i?

OpenStudy (raden):

yes, all the roots exactly : {-5i, 5i, -2, 2}

OpenStudy (anonymous):

ok thank you so much!

OpenStudy (raden):

you're welcome

OpenStudy (anonymous):

could you help me with one more problem? @RadEn

OpenStudy (raden):

i'll try, what is it ?

OpenStudy (anonymous):

find the roots for f(x) = x3 - 5x2 – 25x + 125.

OpenStudy (raden):

x^3 - 5x^2 – 25x + 125 = 0 first, can you find the factors of 125 ?

OpenStudy (anonymous):

1,5,25,125,

OpenStudy (raden):

sometimes we need the negative sign, 125 = {+-1, +-5, +-25, +-125}

OpenStudy (anonymous):

oh ok, so what do we do next?

OpenStudy (raden):

one of those must be as the root of f(x), just trials which they are let it is 5. recheck if this be correct : f(x) = x^3 - 5x^2 – 25x + 125 f(5) = (5)^3 - 5(5)^2 – 25(5) + 125 f(5) = 125 - 125 - 125 + 125 f(5) = 0 because f(5) = 0, so 5 is the root of f(x)

OpenStudy (anonymous):

ok so next we need to find 2 more roots right? because the exponent is 3 for the x

OpenStudy (raden):

yes, that's right. actualy we can use the long division or horner to get the others

OpenStudy (anonymous):

and how would we do that? im sorry im making u do all the work im just trying to learn this

OpenStudy (raden):

nvm, if i can why i didnt help the members here :) see how the diagram horner works :